Question:am i doing this right....
5n/6-1/4=3/8
you times the all the bottm numbers by two
6*2=12
4*2=8
8*2-16
am i doing this right or is there an easyer way

Answer by stanbon(75887)

5n/6-1/4=3/8
-------------
General Rule: When you have an equation with denominators get rid of them.
---------------
The least common multiple of 6, 4 and 8 is 24
Multiply thru the equation by 24 to get:
20n - 6 = 9
20n=15
n=3/4
============
Cheers,
Stan H.